eric and jill plant the garden in 7 hours, working together. jill works 3 times as fast as Eric. In how many hours could jill plant the garden by herself?
Answer by TimothyLamb(4379)

x = eric rate
y = jill rate
---
x + y = 1/7
y = 3x
---
put the system of linear equations into standard form
---
x + y = 0.142857142857
3x - y = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = eric rate = 0.035714286
y = jill rate = 0.10714286
---
r = w/t
t = w/r
t = 1/0.10714286
---
answer:
jill can plant the garden by herself in:
t = 9.333333 hours
t = 9 hours 0.333333*60 minutes
t = 9 hours 20 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php